log(4-3x)-log145=-2
Answer by htmentor(789)

Since log(a/b) = log a - log b, we can combine the logs on the LHS of the equation:
log(4-3x)-log145=-2
log((4-3x)/145) = -2
To "undo" the log, perform the inverse operation:
10^log((4-3x)/145) = 10^(-2)
(4-3x)/145 = 0.01
Solving for x gives:
4-3x = 1.45
x = 0.85
Check:
log(4-3*0.85) - log 145 = log 1.45 - log 145 = -2